# Fully discrete error analysis of finite element discretizations of time-dependent Stokes equations in a stream-function formulation

### v1

@@ 4.1.1 Conformal $C^{1}$ elements
It is natural to seek a finite element space $V_{h}\subset H^{2}_{0}(\Omega)$, i.e. a conformal finite element approximation. In two dimensions, there are a number of such constructions, usually require rather high polynomial order, we will focus only on $C^{1}$ Argyris elements, ( cf. [11, p. 273] for other finite element methods for fourth order problems). Trivially, 1 is satisfied with $\theta=1$ and 2 holds since the method is conformal and the discrete bilinear form is coincides with the continuous one, see Remark 17. In the case of $\Omega$ being convex, the standard duality argument and a priori error estimates show
@@ 4.1.2 The $C^{0}$ interior penalty method
This method is very attractive since it uses Lagrange elements of degree $l\geq 2$. To define the method, we need some additional notation. Let $\mathcal{E}_{h}$ be the set of edges in $\mathcal{T}_{h}$. For $e\in\mathcal{E}_{h}$ and $v\in(X+V_{h})$, with $X$ from Assumption 2, we define the jump $\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]$ of the normal derivative of $v$ across an edge $e$ and the average of the second normal derivative $\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}$ of $v$ on an edge $e$ as follows. If $e\subset\Omega$, we take $n_{e}$ to be one of the two unit vectors normal to $e$. Then $e$ is the common side of two triangles $T_{+}\in\mathcal{T}_{h}$ and $T_{-}\in\mathcal{T}_{h}$, where $n_{e}$ is pointing from $T_{-}$ to $T_{+}$. Thus, on such $e$ we define
@@
$\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]=-\frac{\partial v}{\partial n_{e}}\quad\text{and}\quad\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}=-\frac{\partial^{2}v}{\partial n_{e}^{2}}.$
We define the bilinear form $a_{h}:(X+V_{h})\times(X+V_{h})\to\mathbb{R}$ by
@@
$+\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{(}\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}\left[\hskip-2.52971pt\left[\frac{\partial w}{\partial n}\right]\hskip-2.52971pt\right]+\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}w}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}+\frac{\eta}{|e|}\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]\left[\hskip-2.52971pt\left[\frac{\partial w}{\partial n}\right]\hskip-2.52971pt\right]\Big{)}ds,$
@@
$|e|$ is the length of the edge $e$ and $\eta>0$ is a sufficiently large penalty constant. For this choice of discretization, the results of [10, Lemma 5] show that for any polygonal domain $\Omega$, the 2 holds for $\theta=\frac{1}{2}-\varepsilon$, where $\varepsilon>0$ can be chosen arbitrarily small. It is a standard result, that the 1 is also satisfied. The space $X$ in this case is given by $H^{2}_{0}(\Omega)\cap H^{\frac{5}{2}-\varepsilon}(\Omega)$. Hence, the Ritz projection $R_{h}$ introduced in (45) is well defined for any $\psi\in H^{2}_{0}(\Omega)\cap H^{\frac{5}{2}-\varepsilon}(\Omega)$. Moreover, if $\Omega$ is convex, we have the following error estimate (cf. [10], Theorem 5, with $\alpha=1$)
@@ 5 Numerical results
$\mathbf{f}=\partial_{t}\mathbf{u}-\Delta\mathbf{u}.$
By construction $\mathbf{u}$ satisfies $\nabla\cdot\mathbf{u}=0$, $\mathbf{u}(0)=\mathbf{0}$ and $\mathbf{u}|_{\partial\Omega}=0$, and thus solves the Stokes equations with right hand side $\mathbf{f}$, together with an associated pressure $p=0$. We first discretize the standard weak formulation of these equations by dG(0) in time and the MINI element in space. The numerically observed errors for this setting are reported in Figure 1. We then proceed by modifying the right hand side $\mathbf{f}$ by a gradient field, i.e., set the new right hand side to
$\tilde{\mathbf{f}}=\mathbf{f}+10^{5}\begin{pmatrix}x_{1}^{-0.49}\\ 0\end{pmatrix}$
As $\tilde{\mathbf{f}}$ and $\mathbf{f}$ only differ by a gradient field, they yield the same velocity solution $\mathbf{u}$, and only the pressure changes to $\tilde{p}=p+10^{5}(1.49\cdot x_{1}^{1.49})^{-1}$. As for the discretization of the Stokes with mixed finite elements, the velocity error is bounded by the sum of best approximation errors for velocity and pressure, the error in the velocity increases, which is reported in Figure 2. Lastly we set up a discrete stream-function formulation by taking $g=-\operatorname{curl}(\mathbf{f})$ and solving
$B_{h}(\psi_{kh},\varphi_{kh})=\left(g,\varphi_{kh}\right)_{I\times\Omega}\quad\forall\varphi_{kh}\in X^{r}_{k}(V_{h}).$
We then report the error $\mathbf{u}-\operatorname{\mathbf{curl}}(\psi_{kh})$ in Figure 3. Note that we observe exactly the error estimates derived in Theorem 24. As in the setting of this problem, $\operatorname{curl}$ is applied to the right hand side $\mathbf{f}$ or $\tilde{\mathbf{f}}$ respectively, the same convergence behaviour is obtained for the modified right hand side $\tilde{\mathbf{f}}$, showing the pressure robustness of this method.
Fig. 1: Convergence for discretization with MINI element
Fig. 2: Convergence for discretization with MINI element for Gradient field
Fig. 3: Convergence for Stream-function formulation

### v2

@@ 4.1.1 Conformal $C^{1}$ elements
It is natural to seek a finite element space $V_{h}\subset H^{2}_{0}(\Omega)$, i.e., a conformal finite element approximation. In two dimensions, there are a number of such constructions, usually require rather high polynomial order, we will focus only on $C^{1}$ Argyris elements, (cf. [11, p. 273] for other finite element methods for fourth order problems). Trivially, 1 is satisfied with $\theta=1$ and 2 holds since the method is conformal and the discrete bilinear form is coincides with the continuous one, see Remark 17. In the case of $\Omega$ being convex, the standard duality argument and a priori error estimates show
@@ 4.1.2 The $C^{0}$ interior penalty method
This method is very attractive since it uses Lagrange elements of degree $l\geq 2$. To define the method, we need some additional notation. Let $\mathcal{E}_{h}$ be the set of edges in $\mathcal{T}_{h}$. For $e\in\mathcal{E}_{h}$ and a suitably regular function $v$ we define the jump $\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]$ of the normal derivative of $v$ across an edge $e$ and the average of the second normal derivative $\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}$ of $v$ on an edge $e$ as follows. If $e\subset\Omega$, we take $n_{e}$ to be one of the two unit vectors normal to $e$. Then $e$ is the common side of two triangles $T_{+}\in\mathcal{T}_{h}$ and $T_{-}\in\mathcal{T}_{h}$, where $n_{e}$ is pointing from $T_{-}$ to $T_{+}$. Thus, on such $e$ we define
@@
$\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]=-\frac{\partial v}{\partial n_{e}}\quad\text{and}\quad\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}=\frac{\partial^{2}v}{\partial n_{e}^{2}}.$
It is clear that $\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right],\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}$ are well defined for $v\in V_{h}$. Moreover, following the discussion in Remark 17, the space $X$ from 2 is of the form $X=H^{2}_{0}(\Omega)\cap H^{\frac{5}{2}+\varepsilon}(\Omega)$ for some $\varepsilon>0$. Thus all considered jumps and averages are well defined for $v\in X$. We thus define the bilinear form $a_{h}:(X+V_{h})\times(X+V_{h})\to\mathbb{R}$ by
@@
$+\sum_{e\in\mathcal{E}_{h}}\int_{e}\left(\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}\left[\hskip-2.52971pt\left[\frac{\partial w}{\partial n}\right]\hskip-2.52971pt\right]+\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}w}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}+\frac{\eta}{|e|}\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]\left[\hskip-2.52971pt\left[\frac{\partial w}{\partial n}\right]\hskip-2.52971pt\right]\right)ds,$
@@
$|e|$ is the length of the edge $e$ and $\eta>0$ is a sufficiently large penalty constant. As discussed in Remark 17, for this choice of discretization, the results of [10, Lemma 5] show that for any polygonal domain $\Omega$, the 2 holds for $\theta=\frac{1}{2}-\varepsilon$, where $\varepsilon>0$ can be chosen arbitrarily small. It is a standard result, that the 1 is also satisfied. As mentioned above, the space $X$ in this case is given by $H^{2}_{0}(\Omega)\cap H^{\frac{5}{2}+\varepsilon}(\Omega)$. Hence, the Ritz projection $R_{h}$ introduced in (46) is well defined for any $\psi\in H^{2}_{0}(\Omega)\cap H^{\frac{5}{2}+\varepsilon}(\Omega)$. Moreover, if $\Omega$ is convex, we have the following error estimate (cf. [10], Theorem 5, with $\alpha=1$)
@@ 5 Numerical results
$\mathbf{g}=\partial_{t}\mathbf{u}-\Delta\mathbf{u}.$
By construction $\mathbf{u}$ satisfies $\nabla\cdot\mathbf{u}=0$, $\mathbf{u}(0)=\mathbf{0}$ and $\mathbf{u}|_{\partial\Omega}=0$, and thus solves the Stokes equations with right hand side $\mathbf{g}$, together with an associated pressure $p=0$. We first discretize the standard weak formulation of these equations by dG(0) in time and the MINI element in space. The numerically observed errors for this setting are reported in Figure 1. We then proceed by modifying the right hand side $\mathbf{g}$ by a gradient field, i.e., set the new right hand side to
$\tilde{\mathbf{g}}=\mathbf{g}+10^{5}\begin{pmatrix}x_{1}^{-0.49}\\ 0\end{pmatrix}\in L^{\infty}(I;L^{2}(\Omega)^{2}).$
As $\tilde{\mathbf{g}}$ and $\mathbf{g}$ only differ by a gradient field, they yield the same velocity solution $\mathbf{u}$, and only the pressure changes to $\tilde{p}=p+10^{5}(1.49\cdot x_{1}^{1.49})^{-1}$. As for the discretization of the Stokes with mixed finite elements, the velocity error is bounded by the sum of best approximation errors for velocity and pressure, one can observe in Figure 1, that the error in the velocity increases. Lastly we set up a discrete stream-function formulation by taking $f=-\operatorname{curl}(\mathbf{g})$ and solving
$B_{h}(\psi_{kh},\varphi_{kh})=\left(f,\varphi_{kh}\right)_{I\times\Omega}\quad\text{for all }\varphi_{kh}\in X^{r}_{k}(V_{h}).$
Again we present the error $\mathbf{u}-\operatorname{\mathbf{curl}}(\psi_{kh})$ in Figure 1. Note that we observe exactly the error estimates derived in Theorem 24. Even though the absolute values of the errors are higher than the error for the MINI discretization applied to $\mathbf{g}$, we obtain the same error values, if we applied the stream-function formulation to $\tilde{\mathbf{g}}$, as in the setting of this problem, $\operatorname{curl}$ is applied to the right hand side, and the added gradient field vanishes in the right hand side for $\psi_{kh}$. Thus the same convergence behaviour is obtained for the modified right hand side $\tilde{\mathbf{g}}$, showing the pressure robustness of this method.
$10^{-3}$$10^{-2}$$10^{-1}$$10^{-1.5}$$10^{-1}$$10^{-0.5}$kkMINI ($\mathbf{g}$)MINI ($\tilde{\mathbf{g}}$)streamfct.O(k)O(k)
[…]
Fig. 1: Convergence of discretization with MINI element for different right hand sides (black/red) and interior penalty method for stream-function formulation (blue).